# Abelian surfaces over finite fields containing no curves of genus $3$ or less

### v1

@@ 1. Abelian surfaces containing no curves of genus $\leq 2$
Lemma 1.1 implies that the property of an abelian surface $A$ to contain or not curves of geometric genus $\leq 3$ is an invariant of its isogeny class. In the previous remark, we stressed that this is not the case for curves of arithmetic genus $\leq 2$. We are going to see below that this is also the case when considering curves of arithmetic genus $\leq 3$.
Remark 1.5.
From [14, p. 122], we know that an abelian surface $A$ defined over $\mathbb{F}_{q}$ which is in a not principally polarisable isogeny class splits over the cubic extension of the base field, that is $A\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q^{3}}$ is not simple.
@@
Let $f(t)$ be a Weil polynomial of an abelian surface satisfying one of the equivalent conditions of Theorem 1.2. Then either $f(t)=(t^{2}-q)^{2}$ with $q\in\{2,3\}$ or $f(t)$ is irreducible over $\mathbb{Q}[x]$.
@@
Since $f(t)$ is the Weil polynomial of a simple abelian surface then either it is irreducible or it is the square of a quadratic polynomial. Assume that $f(t)$ is reducible. Then either $f(t)=(t^{2}-q)^{2}$ or $f(t)=(t^{2}-\beta t+q)^{2}$ for some $\beta\in\mathbb{Z}$. If we are the in the former case, then by comparing with Theorem 1.2, we see that $q\in\{2,3\}$, and we are done. So assume that we are in the latter case. Then $a=-2\beta$ and $b=2q+\beta^{2}$. Observe that $b>0$ which implies that $f(t)$ does not satisfy condition (3a). So we must be in case (3b). Hence $a=0$ which implies that $\beta=0$ and $b=2q$, giving a contradiction with the constraints on $b$ from (3b). Therefore $f(t)$ is irreducible. ∎
@@
Let $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$ (resp. $\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$) be the set of Weil polynomials described in Theorem 1.2.(3a) (resp. Theorem 1.2.(3b)) which are irreducible over $\mathbb{Q}[x]$. The set of Weil polynomials described in Theorem 1.2 is partitioned as follows:
@@
Let $A$ be an abelian surface over $\mathbb{F}_{q}$ with Weil polynomial $f(t)=t^{4}+at^{3}+bt^{2}+qat+q^{2}$. Then $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$ if and only if $f(t)$ is irreducible, the Weil polynomial of $A_{\mathbb{F}_{q^{2}}}$ is not irreducible, and $b$ is as in Theorem 1.2.(3b).
@@
It follows from Theorem 1.2.(3b). ∎
@@
Let $A$ be an abelian surface over $\mathbb{F}_{q}$ with Weil polynomial $f(t)=t^{4}+at^{3}+bt^{2}+qat+q^{2}$ in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$. Then $A$ is either ordinary or supersingular. If $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$ then $A$ is ordinary if and only if $a\neq 0$. If $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$, then $A$ is ordinary if and only if $b=1-2q$, or $b=2-2q$ and $p>2$.
@@
Assume that $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$. Note that $A$ cannot have $p$-rank $1$ since in that case it would be principally polarisable by [17, Thm. 4.3] [^1] The authors of [17] call the neither ordinary nor supersingular case “mixed”.. If $a=0$ then we have $b=a^{2}-q=-q$, and therefore $A$ is supersingular. Suppose $a\neq 0$. By [14, Thm. 2], the abelian variety $A$ cannot be supersingular. Hence, $A$ is ordinary. Now, assume that $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$. Obviously if $b=1-2q$ or $b=2-2q$ and $p>2$ then $A$ is ordinary. By Theorem 1.2.(3b), those are the only possible values for $b$ to have ordinarity. To conclude the proof, we need to show that if $A$ is non-ordinary then it is supersingular. Say that $A$ is the Weil restriction of an elliptic curve $E$ over $\mathbb{F}_{q^{2}}$. If $A$ is non-ordinary, then $E$ is supersingular, and hence $A$ is also supersingular. ∎
@@
Let $\alpha,q/\alpha,\beta,q/\beta$ be the complex roots of $f(t)$. Then $-a=\alpha+q/\alpha+\beta+q/\beta$. If $a=0$ then either $\alpha=-\beta$ or $\alpha=-q/\beta$. In both cases $\beta\in\mathbb{Q}(\alpha)$, so the extension is normal and $K$ is Galois. If $a\neq 0$ then $f(t)$ is in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$ and $A$ is ordinary by Lemma 1.9. Hence $K$ is Galois by [12, Lemma 12.1]. ∎
@@ 2. Abelian surfaces containing genus $3$ curves
The key question of the paper is characterising abelian surfaces which do not contain genus $3$ curves. Our main focus will be to characterise such surfaces among those with Weil polynomial in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}\sqcup\{(t^{2}-2)^{2},(t^{2}-3)^{2}\}$, which we know by Theorem 1.2 do not contain curves of geometric (hence arithmetic) genus up to $2$.
We start by analysing the two cases of $(t^{2}-2)^{2}$ and $(t^{2}-3)^{2}$ in Proposition 2.1. Then, in Theorem 2.3, we prove a criterion to check when a simple abelian surface contains a genus $3$ curve. This criterion will be applied in the following sections to characterise which isogeny classes with Weil polynomial in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$ contain an abelian surface containing a curve of genus $3$.
Proposition 2.1.
Let $\mathcal{C}$ (resp. $\mathcal{C}^{\prime}$) be be the isogeny class determined by $(t^{2}-2)^{2}$ (resp. $(t^{2}-3)^{2}$). Then
  • there exists an abelian surface in $\mathcal{C}^{\prime}$ containing an absolutely irreducible smooth curve of genus $3$.
Proof.
Note that there are only $5$ distinct isogeny classes of elliptic curves over $\mathbb{F}_{2}$. A search in the database of isogeny classes abelian threefolds defined over $\mathbb{F}_{2}$ in the LMFDB [16] returns the isogeny classes 3.2.ac_ac_i, 3.2.ab_ac_ e, 3.2.a_ac_a, 3.2.b_ac_ae and 3.2.c_ac_ai. None of them contains a Jacobian, which implies that $A$ does not contain an absolutely irreducible curve of geometric genus $3$. Hence, $A$ does not contain an absolutely irreducible curve of arithmetic genus $3$, as well.
We now apply the same reasoning to $\mathcal{C}^{\prime}$. By searching in the LMFDB, we find $7$ isogeny classes of the form $(t^{2}-3)^{2}f_{E^{\prime}}(t)$ for some elliptic curve $E^{\prime}/\mathbb{F}_{3}$. Moreover, $6$ of these $7$ isogeny classes do not contain a Jacobian. See 3.3.ad_ad_s, 3.3.ac_ad_m, 3.3.ab_ad_g, 3.3.b_ad_ag, 3.3.c_ad_am and 3.3.d_ad_as. The remaining one has label 3.3.a_ad_a and Weil polynomial $(t^{2}-3)^{2}(t^{2}+3)$. It contains only one Jacobian surface, of the curve $C:y^{4}+xz^{3}+2x^{3}z$. The degree $2$ map sending $(x:y:z)\mapsto(x:-y:z)$ defines a double cover of $C$ over the elliptic curve $F\colon y^{2}z+2x^{3}+xz^{2}$. Hence, $C$ is bielliptic. So, by [5, Prop. 1.8], $C$ is contained in an abelian surface in $\mathcal{C}^{\prime}$. ∎
@@
Let $A$ be a simple abelian surface defined over $\mathbb{F}_{q}$. Let $D$ be an effective divisor on $A$ such that $D^{2}=4$ Then, $D$ is an $\mathbb{F}_{q}$-irreducible curve of arithmetic genus $3$. More precisely, $D$ is one of the following:
@@
By the adjunction formula, for an $\mathbb{F}_{q}$-irreducible curve $C$ over $A$ we have $C^{2}=2p_{a}-2$. Thus, the above divisors clearly have self-intersection equal to $4$. Hence, we only need to prove that the list is complete. If $D$ is a single curve then it necessarily has arithmetic genus $3$. If $D=E+F$ is reducible, since it is effective, its components have genus strictly lower than $3$. Since $A$ is simple, the components of $D$ cannot be elliptic curves. Hence they necessarily have arithmetic genus $2$. Then, we get $D^{2}=E^{2}+F^{2}+2\cdot E.F>4$ as $E^{2}=F^{2}=2$ and $E.F>0$ by [5, Lemma 1.1]. ∎
@@
If $A$ does not contain an absolutely irreducible curve of geometric genus $2$, then from Lemma 2.2 we entail that the genus $3$ curve in the statement of Theorem 2.3 is either an absolutely irreducible smooth curve of genus $3$ or an $\mathbb{F}_{q}$-irreducible curve of arithmetic genus $3$ which is not absolutely irreducible.
Having this characterisation at hand, our new goal is to study when an isogeny class contains an abelian surface admitting a polarisation of degree $4$. This question will be answered in Section 4. In light of Proposition 2.1, from now on we will focus on Weil polynomials $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$ only. To start with, we need some technical results concerning the factorisation of $2$ in the extension $K=\mathbb{Q}[t]/f(t)$, which is the main topic of the next section.

### v3

@@ 1. Abelian surfaces containing no curves of genus $\leq 2$
Note that one can distinguish isogeny classes of abelian surfaces admitting no principal polarisation and simple isogeny classes of Weil restrictions by their splitting behavior. From [15, p. 122], we know that an abelian surface $A$ defined over $\mathbb{F}_{q}$ which is in a not principally polarisable isogeny class splits over the cubic extension of the base field, that is $A\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q^{3}}$ is not simple. If $A$ is a simple abelian surface over $\mathbb{F}_{q}$ then $A$ is isogenous to a Weil restriction of an elliptic curve defined over $\mathbb{F}_{q^{2}}$ if and only if $A\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q^{2}}$ is not simple; see [15, Lemma 4].
@@
Let $f(t)$ be a Weil polynomial of an abelian surface satisfying one of the equivalent conditions of Theorem 1.4. Then either $f(t)=(t^{2}-q)^{2}$ with $q\in\{2,3\}$, or $f(t)$ is irreducible over $\mathbb{Q}[x]$.
@@
Since $f(t)$ is the Weil polynomial of a simple abelian surface then either it is irreducible or it is the square of a quadratic polynomial. Assume that $f(t)$ is reducible. Then either $f(t)=(t^{2}-q)^{2}$ or $f(t)=(t^{2}-\beta t+q)^{2}$ for some $\beta\in\mathbb{Z}$. If we are the in the former case, then by comparing with Theorem 1.4, we see that $q\in\{2,3\}$, and we are done. So assume that we are in the latter case. Then $a=-2\beta$ and $b=2q+\beta^{2}$. Observe that $b>0$ which implies that $f(t)$ does not satisfy condition (iii)(iii).(a). So we must be in case (iii)(iii).(b). Hence, $a=0$ which implies that $\beta=0$ and $b=2q$, giving a contradiction with the constraints on $b$ from (iii)(iii).(b). Therefore, $f(t)$ is irreducible. ∎
@@
Let $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$ (resp. $\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$) be the set of Weil polynomials described in Theorem 1.4.(iii)(iii).(a) (resp. Theorem 1.4.(iii)(iii).(b)) which are irreducible over $\mathbb{Q}[x]$. The set of Weil polynomials described in Theorem 1.4 is partitioned as follows:
@@
After having introduced this notation, we are ready to complete the proof of Main Theorem 1.
Corollary 1.10.
Let $f(t)$ be a Weil polynomial defining a simple isogeny class $\mathcal{C}$.
  1. (i)
    $f(t)$ is in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$ if and only if no $A$ in $\mathcal{C}$ contains an $\mathbb{F}_{q}$-irreducible curve with arithmetic genus $\leq 2$.
  2. (ii)
    $f(t)$ is in $\{(t^{2}-2)^{2},(t^{2}-3)^{2}\}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$ if and only if no $A$ in $\mathcal{C}$ contains an absolutely irreducible curve with geometric genus $\leq 2$ and there exists a $B$ in $\mathcal{C}$ which contains an $\mathbb{F}_{q}$-irreducible curve with arithmetic genus $2$.
Proof.
Since $\mathcal{C}$ is simple, no $A$ in $\mathcal{C}$ can contain an $\mathbb{F}_{q}$-irreducible curve of arithmetic genus $1$. Indeed, any such curve would have geometric genus $1$, that is, it would be an elliptic curve which is implies that $A$ is isogenous to the product of two elliptic curves. Hence, Part i follows immediately from Proposition 1.3.
Now, we prove Part ii. Assume first that $f(t)$ is in $\{(t^{2}-2)^{2},(t^{2}-3)^{2}\}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$ and let $A$ be an abelian surface in $\mathcal{C}$. Then, $A$ does not contain absolutely irreducible curves of geometric genus $\leq 2$ by Theorem 1.4. Moreover, there exists an abelian surface in $\mathcal{C}$ which is principally polarised and hence contains an $\mathbb{F}_{q}$-irreducible curve $C$ of arithmetic genus $2$ by Proposition 1.3. Note that this curve $C$ has geometric genus $\leq 2$, so it is not absolutely irreducible because of the assumptions on $A$.
For the converse implication, we first observe that $f(t)$ is in $\{(t^{2}-2)^{2},(t^{2}-3)^{2}\}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$ by Theorem 1.4. By Proposition 1.3, we have that $B$ is principally polarised and thus $f(t)\not\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$. ∎
@@
Let $A$ be an abelian surface over $\mathbb{F}_{q}$ with Weil polynomial $f(t)=t^{4}+at^{3}+bt^{2}+qat+q^{2}$. Then $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$ if and only if $f(t)$ is irreducible, the Weil polynomial of $A_{\mathbb{F}_{q^{2}}}$ is not irreducible, and $b$ is as in Theorem 1.4.(iii)(iii).(b).
@@
It follows from Theorem 1.4.(iii)(iii).(b) and Remark 1.7. ∎
@@
Let $A$ be an abelian surface over $\mathbb{F}_{q}$ with Weil polynomial $f(t)=t^{4}+at^{3}+bt^{2}+qat+q^{2}$ in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$. Then $A$ is either ordinary or supersingular. If $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$ then $A$ is ordinary if and only if $a\neq 0$. If $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$, then $A$ is ordinary if and only if $b=1-2q$ or, $b=2-2q$ and $p>2$.
@@
Assume that $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$. Note that $A$ cannot have $p$-rank $1$ since in that case it would be principally polarisable by [18, Thm. 4.3] [^2] The authors of [18] call the neither ordinary nor supersingular case “mixed”.. If $a=0$ then we have $b=a^{2}-q=-q$, and therefore $A$ is supersingular. Suppose $a\neq 0$. By [15, Thm. 2], the abelian variety $A$ cannot be supersingular. Hence, $A$ is ordinary. Now, assume that $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$. Obviously if $b=1-2q$, or $b=2-2q$ and $p>2$ then $A$ is ordinary. By Theorem 1.4.(iii)(iii).(b), those are the only possible values for $b$ to have ordinarity. To conclude the proof, we need to show that if $A$ is non-ordinary then it is supersingular. Say that $A$ is the Weil restriction of an elliptic curve $E$ over $\mathbb{F}_{q^{2}}$. If $A$ is non-ordinary, then $E$ is supersingular, and hence $A$ is also supersingular. ∎
@@
Let $\alpha,q/\alpha,\beta,q/\beta$ be the complex roots of $f(t)$. Then $-a=\alpha+q/\alpha+\beta+q/\beta$. If $a=0$ then either $\alpha=-\beta$ or $\alpha=-q/\beta$. In both cases $\beta\in\mathbb{Q}(\alpha)$, so the extension is normal and $K$ is Galois. If $a\neq 0$ then $f(t)$ is in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$ and $A$ is ordinary by Lemma 1.12. Hence $K$ is Galois by [13, Lemma 12.1]. ∎
@@ 2. Abelian surfaces containing genus $3$ curves
The key question of the paper is characterising abelian surfaces which do not contain genus $3$ curves. After a technical lemma, we prove Main Theorem 2, stated below as Theorem 2.2.
@@
Let $A$ be a simple abelian surface defined over $\mathbb{F}_{q}$. Let $D$ be an effective rational divisor on $A$ such that $D^{2}=4$. Then, $D$ is an $\mathbb{F}_{q}$-irreducible curve of arithmetic genus $3$. More precisely, $D$ is one of the following:
@@
By the adjunction formula, for an $\mathbb{F}_{q}$-irreducible curve $D$ over $A$ we have $D^{2}=2p_{a}(D)-2$. If $D$ is an $\mathbb{F}_{q}$-irreducible curve then it necessarily has arithmetic genus $3$, and it must be one of the three possibilities listed in the statement. Hence, to conclude the proof it suffices to show that $D$ is an $\mathbb{F}_{q}$-irreducible curve.
If $D=E+F$ were reducible, using Equation (2) and the assumption that $D^{2}=4$, we would get that it has at most $2$ irreducible components having arithmetic genus strictly lower than $3$. Since $A$ is simple, the components $E$ and $F$ of $D$ cannot be elliptic curves. Hence, they necessarily have arithmetic genus $2$. So, again by the adjunction formula, we get that $E^{2}=F^{2}=2$. Moreover, we have $E.F>0$ by Lemma 1.2. Combining these with Equation (2), we get $D^{2}>4$ which is a contradiction. It follows that $D$ must be an $\mathbb{F}_{q}$-irreducible curve, thus concluding the proof. ∎
@@
If $A$ does not contain an absolutely irreducible curve of geometric genus $2$, then, from Lemma 2.1, we deduce that the genus $3$ $\mathbb{F}_{q}$-irreducible curve in the statement of Theorem 2.2 is either an absolutely irreducible smooth curve of genus $3$ or an $\mathbb{F}_{q}$-irreducible curve of arithmetic genus $3$ which is not absolutely irreducible.
From now on, our main focus will be to characterise the abelian surfaces with a polarisation of degree $4$ among those with Weil polynomial in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}\sqcup\{(t^{2}-2)^{2},(t^{2}-3)^{2}\}$, which as we know by Theorem 1.4 do not contain curves of geometric (hence arithmetic) genus up to $2$.
We start by analysing the two cases of $(t^{2}-2)^{2}$ and $(t^{2}-3)^{2}$ in Proposition 2.4. In the following sections, we will apply Theorem 2.2 to characterise which isogeny classes with Weil polynomial in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$ contain an abelian surface containing a curve of genus $3$.
Proposition 2.4.
Let $\mathcal{C}$ (resp. $\mathcal{C}^{\prime}$) be the isogeny class determined by $(t^{2}-2)^{2}$ (resp. $(t^{2}-3)^{2}$). Then
  • there exists an abelian surface in $\mathcal{C}^{\prime}$ containing the absolutely irreducible smooth curve of genus $3$ defined by $y^{4}+2x^{3}z+xz^{3}$.
Proof.
Note that there are only $5$ distinct isogeny classes of elliptic curves over $\mathbb{F}_{2}$. A search in the database of isogeny classes of abelian threefolds defined over $\mathbb{F}_{2}$ in the LMFDB [17] returns the isogeny classes 3.2.ac_ac_i, 3.2.ab_ac_ e, 3.2.a_ac_a, 3.2.b_ac_ae and 3.2.c_ac_ai. None of them contains a Jacobian, which implies that $A$ does not contain an absolutely irreducible curve of geometric genus $3$. Hence, $A$ does not contain an absolutely irreducible curve of arithmetic genus $3$, as well.
We now apply a similar reasoning to $\mathcal{C}^{\prime}$. By searching in the LMFDB, we find $7$ isogeny classes of the form $(t^{2}-3)^{2}f_{E^{\prime}}(t)$ for some elliptic curve $E^{\prime}/\mathbb{F}_{3}$. Moreover, $6$ of these $7$ isogeny classes do not contain a Jacobian. See 3.3.ad_ad_s, 3.3.ac_ad_m, 3.3.ab_ad_g, 3.3.b_ad_ag, 3.3.c_ad_am and 3.3.d_ad_as. The remaining one has label 3.3.a_ad_a and Weil polynomial $(t^{2}-3)^{2}(t^{2}+3)$. It contains only the Jacobian surface $J$ of the curve $C:y^{4}+2x^{3}z+xz^{3}$. The degree $2$ involution sending $(x:y:z)\mapsto(x:-y:z)$ defines a double cover $\pi:C\to F$ where $F$ is the elliptic curve given by $y^{2}z+2x^{3}+xz^{2}$. Hence, $C$ is bielliptic, and $J/\pi^{*}F\in\mathcal{C}^{\prime}$. The branch points of $\pi$ are the images of the fixed points of the involution, that is, the points on $C$ with $y=0$. We get the points $(0:0:1)$, $(1:0:1)$ and $(2:0:1)$ together with the point at infinity $(1:0:0)$. We now follow Barth’s proof [5, Prop. 1.8] (the forward direction of the implication), where the author defines a map $\varphi:C\hookrightarrow J\to J/\pi^{*}F$ using any of the branch points of $\pi$. A careful analysis of the construction shows that, since the branch point is defined over $\mathbb{F}_{3}$, the same holds for $\varphi$. Barth’s argument to show that $\varphi$ is an embedding carries over to finite fields without any modification. This concludes the proof. ∎
Having this characterisation at hand, we can easily derive the Corollary in the introduction.
Corollary 2.5.
Let $f(t)$ be the Weil polynomial of a simple isogeny class $\mathcal{C}$ of abelian surfaces over $\mathbb{F}_{q}$.
  1. (i)
    Assume that $f(t)\in\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}\sqcup\mathcal{P}^{\mathrm{irr}}_{\mathrm{Wres}}$ and let $A$ be an abelian surface in $\mathcal{C}$. If $A$ contains an absolutely irreducible curve of arithmetic genus $\leq 3$ then $A$ admits a polarisation of degree $4$.
  2. (ii)
    No abelian surface in $\mathcal{C}$ contains an $\mathbb{F}_{q}$-irreducible curve of arithmetic genus $\leq 3$ if and only if $f(t)$ is in $\mathcal{P}^{\mathrm{irr}}_{\mathrm{npp}}$ and no abelian surface in $\mathcal{C}$ admits a polarisations of degree $4$.
Proof.
We start by proving i. Let $C$ be an absolutely irreducible curve of arithmetic genus $\leq 3$ on $A$. By the assumption on $f(t)$, Theorem 1.4 implies that $p_{a}(C)=3$. Therefore, $A$ admits a polarisation of degree $4$ by Theorem 2.2.
We now continue with ii. Since $\mathcal{C}$ is simple, no $A$ in $\mathcal{C}$ contains an $\mathbb{F}_{q}$-irreducible curve of arithmetic or geometric genus $1$, because such a curve would be an elliptic curve and hence an isogeny factor of $\mathcal{C}$. Then we conclude the proof by observing that Proposition 1.3 and Theorem 2.2 imply that an abelian surface $A$ in $\mathcal{C}$ contains an $\mathbb{F}_{q}$-irreducible curve of arithmetic genus $\leq 3$ if and only if $A$ admits a polarisation which is of degree $1$ or $4$. ∎